1. Write down the definition of the derivative and use it to find f(x) for f(x)=5x² - 2x
Answer #1

* Given f (x) = 5x² - 2x diff worot & f(x) = a (3x² - 2x Apply linearity 5 d (x) - 2 d dx (x) dx ad id (a)=nxhi = 5 5(2x2-